Consider the reaction shown.



4HCl(g)+O2(g)⟶2Cl2(g)+2H2O(g)



Calculate the number of grams of Cl2 formed when 0.225 mol HCl reacts with an excess of O2 .

1
Expert's answer
2022-03-19T08:52:35-0400

n(HCl) = 0.225 mol

m(Cl2) - ?

Solution:

n(Cl2) = n(HCl)/2 = 0.225 mol/2 = 0.113 mol

m(Cl2) = n(Cl2)*M(Cl2) = 0.113mol*71g/mol

= 8.02 g

Answer: mass of Cl2 8.02 g